# A Financial Brain Scan of the LLM

### v1

@@ A Interpretable LLMs’ embeddings
Having shown that efficiency is not compromised and that all features contribute to maximum performance, we now turn to the main advantage of sparse representations: interpretability. Each feature in a sparse representation corresponds to a specific semantic meaning. In this paper, we use the pre-trained and labeled sparse model released by Google DeepMind (Lieberum et al., 2024). [^4] The features in this model are labeled using the methodology described in Section III. The features’ labels are not finance-specific but reflect general applications. Some labels capture subtle distinctions between related concepts. For example, one feature is labeled “negative sentiments conditions acceptance limitations,” while another is labeled “expressing opinions judgments performance events.” Although these can be viewed as distinct categories, it may be more appropriate to group them under broader “sentiment” categories. Thus, constructing groups of labels is a natural step.
@@ B Steering and Bias Correction
We start by building a long-short trading strategy based on prompt-extracted sentiment from news in the spirit of Lopez-Lira and Tang (2023a) and Chen et al. (2024). We employ an LLM to estimate the sentiment of aftermarket and overnight Reuters news concerning individual stocks. On the sparse representation, we select a feature linked to “positivity’. For each news item, we obtain a baseline LLM forecast and additional forecasts corresponding to varying levels of positive steering. As expected, the proportion of positive classifications increases monotonically with stronger positive steering and the conditional returns are consistent with the model’s induced interpretation of the same headline. In the spirit of Lopez-Lira and Tang (2023a), we use these forecasts to construct intra-day long–short portfolios, going long (equally weighted) on stocks with positive sentiment and short on those with negative sentiment. Interestingly, negatively steered predictions achieve a statistically significant higher annualized Sharpe ratio (4.28) compared to the baseline (3.87).666 We follow the approach proposed by Jobson and Korkie (1981), incorporating the correction noted in Memmel (2003), to test the significance of our results. In addition, we show that the negatively steered portfolio generates statistically significant alpha relative to the baseline strategy. Notably, all negatively steered predictions yield higher Sharpe ratios than both the baseline and positively steered predictions, with the effect most pronounced for moderate steering.
@@ E Concept Steering
The first benefit of SAE-augmented LLMs, discussed in Section V, is the sparse representation itself, which enables the construction of interpretable text embeddings. These embeddings help understand how LLMs process financial documents. The second benefit is that labeled inner representations facilitate the possibility of pushing or steering LLMs toward specific concepts such as risk aversion, sentiment polarity, or other preferences and beliefs (Section VI).
@@
Assume a SAE has been trained on the $m^{th}$ residual stream, with encoder121212 In practice, SAE can be trained on any intermediary layer, but for the sake of clarity we focus on this simplified example. The same logic applies if the perturbation occurs earlier in the LLMs.
@@
As we show in Section VI.B, if $z_{j}^{(m)}$ is associated with positive sentiment, texts generated using $P^{\text{(Steered)}}$ are likely to be more optimistic than those generated without steering. Conversely, steering toward risk-related features produces more cautious-sounding text.
@@ IV Data
Second, our analyses are either comparative—demonstrating that sparse embeddings outperform alternative embeddings derived from the same LLM (Sections V.A and VI)—or centered on interpretability rather than predictive performance (Section V.B).
@@
V Interpreting the Embedding
In this section, we examine the economic performance and interpretability of the sparse embeddings produced by the SAEs (see Section III). Unlike simpler bag-of-words embeddings (see, e.g., Tetlock, 2007), SAE embeddings are derived from a pre-trained LLM and therefore capture subtle textual nuances that can be informative in a wide range of applications. Moreover, in contrast to previously proposed LLM-based embeddings (Sarkar, 2025), SAE embeddings are explicitly designed to be interpretable.
@@ A Performance
    We use Gemma9B with augmented SAEs to generate sparse, interpretable vectors ($\nu_{i,t}$) for each news article in our sample (see Section IV).
@@
Recall that SAE embeddings are sparse and semantically meaningful representations of an LLM’s text processing, in our case Gemma9B. Hence, this notion of a “virtue of complexity” extends to LLMs analyzing financial news more broadly. Indeed, the exercise in this section and the prompting-based sentiment estimation approach latter discussed in Section VI or Lopez-Lira and Tang (2023a); Chen et al. (2024) are closely related. In both cases, we employ part of an LLM to compress news into a lower-dimensional representation. In the prompting approach, this representation is further processed by the LLM to generate text from which predictions are parsed. In the approach presented here, the lower-dimensional representation is instead mapped directly into an interpretable embedding, which is then used to predict returns.
@@ B Interpretability
Having shown in the previous section that our embedding does not sacrifice performance, we now turn to its main advantage over previous methods: interpretability. Each feature in our sparse embedding is labeled through the following procedure: (1) a large corpus of texts is processed by the LLM, (2) texts that “activate” a given feature (i.e., yield a non-zero value) are grouped together, and (3) another LLM is prompted to analyze these texts and assign an appropriate label. A detailed description of these steps is provided in Section III.
@@
Our experiment in Section V.A, which uses 5,000 features as input and produces an out-of-sample score of 5.51, can therefore be interpreted as being mapped to 5,000 distinct feature labels. Since this number is too large for direct interpretation, we construct higher-level clusters of features using the following steps:
@@
The second cluster of interest is Punctuation and Symbols. The corresponding wordcloud, shown in panel (g) of Figure 4, displays words more strongly associated with coding and programming than with punctuation per se. This peculiarity arises from the feature labeling procedure. As described in Section III, labels are assigned by passing a large general-purpose textual dataset through the LLM and recording the types of texts or lines most frequently activating a given feature. Since a large share of code consists of punctuation (for example, most C++ lines end with “;” and colons initiate conditional or loop statements in Python), the labeling process mechanically associates the cluster with punctuation. We confirmed this interpretation by manually inspecting the texts linked to feature labeling.
@@
We use these clusters to estimate prediction models based solely on the features within each cluster. In other words, we replicate the procedure described in Section V.A, but instead of using all 5,000 features, we restrict the input to the subset belonging to a given cluster. This exercise produces 17 sets of out-of-sample forecasts, one for each cluster. Figure 5 reports the pairwise correlations among these forecasts.
@@
We next construct an alternative set of predictions in the spirit of Shapley values (Gu et al., 2020). For each labeled cluster (e.g., Sentiment, Risks), we form a feature set consisting of all 5,000 features excluding those belonging to that cluster. We then apply the procedure described in Section V.A to obtain prediction accuracies and the Sharpe ratios of the corresponding long–short strategies.
@@
VI Steering LLM Sentiment
We now turn to applications of steering, discussed in detail in Section III.E. In essence, the approach consists of identifying relevant concepts in the SAE’s representation (e.g., risk aversion or optimism/pessimism) and selectively amplifying or attenuating them during text generation, thereby producing outputs steered toward those concepts. This approach is fundamentally different from prompt engineering.
@@ VI Steering LLM Sentiment
Before turning to the next section, where we examine how steering can be applied to de-bias LLMs in financial contexts, we first present a simple “proof of concept” to test whether steering produces the expected outcome in an economic classification task. Using the news dataset described in Section IV, we prompt the LLM to classify each news item as positive or negative with the following instruction:
@@
We then repeat the procedure while steering the model: by intervening on a node linked to positivity/optimism, we steer the node in either the positive or negative direction. [^16] Index 111712 (expressions of positive sentiment and appreciation) on google/gemma-scope-9b-it-res (layer 20) with 131k features. This generates a set of predictions in which, for each news item, we obtain classifications under a grid of positive and negative steering levels, alongside a benchmark with no steering. As in Section V, we use the next day’s intraday return as the classification label; that is, each news item is paired with the corresponding firm’s realized return.
@@ A Optimism Bias and LLMs
These findings validate steering as a tool to both identify and mitigate LLM bias. By selecting an appropriate steering level for a given feature, one can correct biases such as those documented in Fedyk et al. (2024). Conversely, steering can also be employed to deliberately amplify bias for research purposes—for instance, to simulate varying levels of risk aversion or home bias in human decision-making. We discuss this broader methodological potential in Section VI.B.
@@ A Topic Merging
Section V.B describes the use of the k-means algorithm (McQueen, 1967) to cluster label features into 25 distinct groups. The choice of 25 clusters is guided by the maximization of the silhouette coefficient (Rousseeuw, 1987), consistent with standard practice in the literature.
@@
As discussed in Section V.B, when clusters were difficult to interpret, we consulted the documentation available at https://www.neuronpedia.org to identify the types of texts used to generate the reported labels. For instance, although the labels in Topic 0 appear related to programming, inspection of the source texts revealed that this association arises because punctuation symbols—frequently occurring in code—were prominent in the labeling process.

### v2

@@ A Interpretable LLMs’ embeddings
Having shown that efficiency is not compromised and that all features contribute to maximum performance, we now turn to the main advantage of sparse representations: interpretability. Each feature in a sparse representation corresponds to a specific semantic meaning. In this paper, we use the pre-trained and labeled sparse model released by Google DeepMind (Lieberum et al., 2024). [^3] The features in this model are labeled using the methodology described in Section II. The features’ labels are not finance-specific but reflect general applications. Some labels capture subtle distinctions between related concepts. For example, one feature is labeled “negative sentiments conditions acceptance limitations,” while another is labeled “expressing opinions judgments performance events.” Although these can be viewed as distinct categories, it may be more appropriate to group them under broader “sentiment” categories. Thus, constructing groups of labels is a natural step.
@@ B Steering and Bias Correction
We start by building a long-short trading strategy based on prompt-extracted sentiment from news in the spirit of Lopez-Lira and Tang (2023a) and Chen et al. (2024). We employ an LLM to estimate the sentiment of aftermarket and overnight Reuters news concerning individual stocks. On the sparse representation, we select a feature linked to “positivity’. For each news item, we obtain a baseline LLM forecast and additional forecasts corresponding to varying levels of positive steering. As expected, the proportion of positive classifications increases monotonically with stronger positive steering and the conditional returns are consistent with the model’s induced interpretation of the same headline. In the spirit of Lopez-Lira and Tang (2023a), we use these forecasts to construct intra-day long–short portfolios, going long (equally weighted) on stocks with positive sentiment and short on those with negative sentiment. Interestingly, negatively steered predictions achieve a statistically significant higher annualized Sharpe ratio (4.28) compared to the baseline (3.87). [^5] We follow the approach proposed by Jobson and Korkie (1981), incorporating the correction noted in Memmel (2003), to test the significance of our results. In addition, we show that the negatively steered portfolio generates statistically significant alpha relative to the baseline strategy. Notably, all negatively steered predictions yield higher Sharpe ratios than both the baseline and positively steered predictions, with the effect most pronounced for moderate steering.
@@ E Concept Steering
The first benefit of SAE-augmented LLMs, discussed in Section IV, is the sparse representation itself, which enables the construction of interpretable text embeddings. These embeddings help understand how LLMs process financial documents. The second benefit is that labeled inner representations facilitate the possibility of pushing or steering LLMs toward specific concepts such as risk aversion, sentiment polarity, or other preferences and beliefs (Section V).
@@
Assume a SAE has been trained on the $m^{th}$ residual stream, with encoder [^11] In practice, SAE can be trained on any intermediary layer, but for the sake of clarity we focus on this simplified example. The same logic applies if the perturbation occurs earlier in the LLMs.
@@
As we show in Section V.B, if $z_{j}^{(m)}$ is associated with positive sentiment, texts generated using $P^{\text{(Steered)}}$ are likely to be more optimistic than those generated without steering. Conversely, steering toward risk-related features produces more cautious-sounding text.
@@ III Data
Second, our analyses are either comparative—demonstrating that sparse embeddings outperform alternative embeddings derived from the same LLM (Sections IV.A and V)—or centered on interpretability rather than predictive performance (Section IV.B).
@@
IV Interpreting the Embedding
In this section, we examine the economic performance and interpretability of the sparse embeddings produced by the SAEs (see Section II). Unlike simpler bag-of-words embeddings (see, e.g., Tetlock, 2007), SAE embeddings are derived from a pre-trained LLM and therefore capture subtle textual nuances that can be informative in a wide range of applications. Moreover, in contrast to previously proposed LLM-based embeddings (Sarkar, 2025), SAE embeddings are explicitly designed to be interpretable.
@@ A Performance
    We use Gemma9B with augmented SAEs to generate sparse, interpretable vectors ($\nu_{i,t}$) for each news article in our sample (see Section III).
@@
Recall that SAE embeddings are sparse and semantically meaningful representations of an LLM’s text processing, in our case Gemma9B. Hence, this notion of a “virtue of complexity” extends to LLMs analyzing financial news more broadly. Indeed, the exercise in this section and the prompting-based sentiment estimation approach latter discussed in Section V or Lopez-Lira and Tang (2023a); Chen et al. (2024) are closely related. In both cases, we employ part of an LLM to compress news into a lower-dimensional representation. In the prompting approach, this representation is further processed by the LLM to generate text from which predictions are parsed. In the approach presented here, the lower-dimensional representation is instead mapped directly into an interpretable embedding, which is then used to predict returns.
@@ B Interpretability
Having shown in the previous section that our embedding does not sacrifice performance, we now turn to its main advantage over previous methods: interpretability. Each feature in our sparse embedding is labeled through the following procedure: (1) a large corpus of texts is processed by the LLM, (2) texts that “activate” a given feature (i.e., yield a non-zero value) are grouped together, and (3) another LLM is prompted to analyze these texts and assign an appropriate label. A detailed description of these steps is provided in Section II.
@@
Our experiment in Section IV.A, which uses 5,000 features as input and produces an out-of-sample score of 5.51, can therefore be interpreted as being mapped to 5,000 distinct feature labels. Since this number is too large for direct interpretation, we construct higher-level clusters of features using the following steps:
@@
The second cluster of interest is Punctuation and Symbols. The corresponding wordcloud, shown in panel (g) of Figure 4, displays words more strongly associated with coding and programming than with punctuation per se. This peculiarity arises from the feature labeling procedure. As described in Section II, labels are assigned by passing a large general-purpose textual dataset through the LLM and recording the types of texts or lines most frequently activating a given feature. Since a large share of code consists of punctuation (for example, most C++ lines end with “;” and colons initiate conditional or loop statements in Python), the labeling process mechanically associates the cluster with punctuation. We confirmed this interpretation by manually inspecting the texts linked to feature labeling.
@@
We use these clusters to estimate prediction models based solely on the features within each cluster. In other words, we replicate the procedure described in Section IV.A, but instead of using all 5,000 features, we restrict the input to the subset belonging to a given cluster. This exercise produces 17 sets of out-of-sample forecasts, one for each cluster. Figure 5 reports the pairwise correlations among these forecasts.
@@
We next construct an alternative set of predictions in the spirit of Shapley values (Gu et al., 2020). For each labeled cluster (e.g., Sentiment, Risks), we form a feature set consisting of all 5,000 features excluding those belonging to that cluster. We then apply the procedure described in Section IV.A to obtain prediction accuracies and the Sharpe ratios of the corresponding long–short strategies.
@@
V Steering LLM Sentiment
We now turn to applications of steering, discussed in detail in Section II.E. In essence, the approach consists of identifying relevant concepts in the SAE’s representation (e.g., risk aversion or optimism/pessimism) and selectively amplifying or attenuating them during text generation, thereby producing outputs steered toward those concepts. This approach is fundamentally different from prompt engineering.
@@ V Steering LLM Sentiment
Before turning to the next section, where we examine how steering can be applied to de-bias LLMs in financial contexts, we first present a simple “proof of concept” to test whether steering produces the expected outcome in an economic classification task. Using the news dataset described in Section III, we prompt the LLM to classify each news item as positive or negative with the following instruction:
@@
We then repeat the procedure while steering the model: by intervening on a node linked to positivity/optimism, we steer the node in either the positive or negative direction. [^15] Index 111712 (expressions of positive sentiment and appreciation) on google/gemma-scope-9b-it-res (layer 20) with 131k features. This generates a set of predictions in which, for each news item, we obtain classifications under a grid of positive and negative steering levels, alongside a benchmark with no steering. As in Section IV, we use the next day’s intraday return as the classification label; that is, each news item is paired with the corresponding firm’s realized return.
@@ A Optimism Bias and LLMs
These findings validate steering as a tool to both identify and mitigate LLM bias. By selecting an appropriate steering level for a given feature, one can correct biases such as those documented in Fedyk et al. (2024). Conversely, steering can also be employed to deliberately amplify bias for research purposes—for instance, to simulate varying levels of risk aversion or home bias in human decision-making. We discuss this broader methodological potential in Section V.B.
@@ A Topic Merging
Section IV.B describes the use of the k-means algorithm (McQueen, 1967) to cluster label features into 25 distinct groups. The choice of 25 clusters is guided by the maximization of the silhouette coefficient (Rousseeuw, 1987), consistent with standard practice in the literature.
@@
As discussed in Section IV.B, when clusters were difficult to interpret, we consulted the documentation available at https://www.neuronpedia.org to identify the types of texts used to generate the reported labels. For instance, although the labels in Topic 0 appear related to programming, inspection of the source texts revealed that this association arises because punctuation symbols—frequently occurring in code—were prominent in the labeling process.